A man deposited P2,000 in a bank at a rate at 15% per annum from March 21, 1996 to October 25, 1997. Find the exact simple interest.?
Answer by Theo(13342)

to solve this exactly, you would need to know the number of days in each month involved in 1996 and 1997.

since excel has a calendar function, i used that to determine the number of days between march 21, 1996 and october 25, 1997.

excel told me that the number of days between march 21, 1996 and october 25, 1997 is equal to 522.

the answer to this question involves how banks calculate daily interest rate.

if they assume 365 days in a year, then the daily interest rate would be .15 / 365 = 4.109589041 * 10^-4.

if they assume 360 days in a year, then the then the daily interest rate would be .15 / 360 = 4.1666666667 * 10^-4.

it all depends on what the banks assume, which is what makes calculating the exact simple interest difficult in this problem.

you just need to know what the banks assume when they are determining the daily interest rate.


bottom line:

from march 21, 1996 to august 25, 1997, excel calculates 522 days.

if the daily interest rate is calculated based on 365 days in a year, then the daily interest rate will be .15 / 365 = 4.109589041 * 10^-4.

if the daily interest rate is calculated based on 360 days in a year, then the daily interest rate will be .15 / 360 = 4.1666666667 * 10^-4.

the formula for calculating simple interest if i = p * r * n.

i is the interest.
p is the principal.
r is the interest rate per time period.
n is the number of time periods.

since you are dealing with daily time periods, then the formula becomes:

if the daily interest rate assumes 365 days, then i = 2000 * 4.109589041 * 10^-4 * 522 = 429.0410959.

if the daily interest rate assumes 360 days, then i = 2000 * 4.1666666667 * 10^-4 * 522 = 435.